Make New Entry,
Make Followup Entry
User name brads
Log entry time 22:01:07 on March 10, 2012
Entry number 365512
keyword=GEp/G2p status update
We had a problem with the RHRS dipole controller being 'stuck' starting
last night. We continued with BCM, BPM, and a limited optics study until
Saturday morning. Jack came in Saturday morning, we took a Controlled
Access, and he was able to reestablish control, ramp the dipole down and
reset the controller. BCM calibrations continued while we waited for the
dipole to ramp back up to the production setpoint.
The target group started preparation for production running around 2pm
and we had a polarized target around 5:30--6pm. The optics matrix for
the RHRS is not fully optimized, but is sufficient to take production
data for GEp. After consulting with Doug we decided to postpone the
remain optics study and get some data to disk. GEp is now taking its
first production data.
Issues to watch for (these are also on the whiteboard):
- Shift staffing is a real problem! There are many vacancies that
must be filled ASAP. Come on folks, step up and sign up!
- Beam leakage from Hall C can be significant, shift crew are
instructed to verify that the Hall A Laser slit is set to 14. They
should also be requesting a leakage check every 4--5 hours just to be safe.
- The IHWP will be flipped at Hall C's request every 4 hours starting
this evening. Hall C and/or MCC should inform us when this happens so we
can log the change and stop/start our runs. Hall C would like to shorten
this flip interval in the future -- we'll negotiate this
on Sunday or Monday.
- The target magnet current spontaneously drops by ~0.5A occasionally
(look in Halog for "Target Current" entries). The Target operator needs
to watch for these and know how to correct the issue -- ask the TO how to
do this when you come on shift if you don't already know.
- Disable some trigger plane bars in the R/LHRSs to optimize trigger
(ie. cut out some of the Nitrogen quasi-elastic triggers). We want to do
this carefully -- a conservative cut may happen tonight, with further
- Finish optics data taking. This is a G2p priority -- we will get
back to this Sunday/Monday.
- HARP 1H05A still isn't working.
- Complete BCM/Tungsten Calorimeter calibration.
A copy of this log entry has been emailed to: jpchen,slifer,camsonne,doug,rom,brads,g2p